OpenStudy (anonymous):

Think i finally got this:: |-3|^2 + 5^2 - 4^2 is 36! Please tell me i got this right :))

OpenStudy (anonymous):

9+25-16=36?

OpenStudy (anonymous):

i got 27+25-16=36 :/

OpenStudy (anonymous):

Ah, looks like you did 3^3 instead of 3^2.

OpenStudy (phi):

|-3| is 3 and 3^2 is 3*3 = 9

OpenStudy (anonymous):

sorry its suppose to be -3^3

OpenStudy (anonymous):

|-3|^3 + 5^2 - 4^2 = 36

OpenStudy (anonymous):

Oh, then it's all good.

OpenStudy (anonymous):

thx :)